OpenStudy (jmartinez638):

Find all complex solution to the equation z^6 = 64.

OpenStudy (jmartinez638):

Any help appreciated!

OpenStudy (snowsurf):

\[z = \sqrt[6]{64}\]

OpenStudy (snowsurf):

So z = 2. There two possibilities which is 2 ad -2. -2i would be the complex root if 64 was negative.

OpenStudy (jmartinez638):

So is the only possible complex root of the equation -2i, or would there be no complex root with the equation in its current configuration?

OpenStudy (snowsurf):

Yes. Complex roots only arise when the value is negative in the square root sign. So technically this equation only has real roots. If this equation was \[z =\sqrt[6]{-64}\] then you would have -2i and 2i as your roots. Since 64 is positive you would have 2 and -2.

OpenStudy (loser66):

It is not enough to find all the complex roots out. You have to go some more steps.

OpenStudy (loser66):

Or, I am making it so complicated. :)

OpenStudy (snowsurf):

That is about it. No extra work is needed.

OpenStudy (loser66):

Like this, if z=2, then \(z=|2|e^{i*0}\) Hence there exists a \(w\) as a root of z such that \(z^6 =64\) Let \(w=|w|e^{i\tau}\) \(w^6=|w|^6 e^{6i\tau}\) Hence \(|w| =2\) Now, imaginary part \(6i\tau =i*0 +2i kpi \) hence \(\tau = k\pi/3, 0\leq k\leq 5\) for k =0, you have \(w=2\) for k =1, you have \(w=2e^{i\pi/3}=2(cos(\pi/3)+isin(\pi/3)=1+i\sqrt 3\) Now, check up \((1+i\sqrt 3)^6=64\) Hence, one of value of z is \(1+i\sqrt 3\) Do the same with k =2,3,4,5 to find all of complex roots of z

OpenStudy (loser66):

This is the proof of \((1+i\sqrt 3)^6=64\) https://www.wolframalpha.com/input/?i=(1%2Bisqrt(3))%5E6

OpenStudy (loser66):

@snowsurf I think I complicated the stuff. hehehe....

OpenStudy (snowsurf):

Haha. That is a cool result. I would never had thought of that.

OpenStudy (loser66):

:) it is from my complex analysis course.

OpenStudy (snowsurf):

So you applied Euler's formula

OpenStudy (anonymous):

\[z^6=64=r \left( \cos \theta+\iota \sin \theta \right)\] \[r \cos \theta=64,r \sin \theta=0,\] square and add \[r^2\left( \cos ^2 \theta +\sin ^2\theta \right)=64^2+0^2,r=64\] Hence \[\cos \theta=1=\cos 0=\cos( 0+2 n \pi),\theta=2n \pi\] where n is an integer. \[z^6=64(\cos 2 n \pi+\iota \sin 2n \pi)=64e ^{2n \pi}\] \[z=64^{\frac{ 1 }{ 6 }}e ^{\frac{ 2n \pi }{ 6 }}=2e ^{\frac{ n \pi }{ 3 }}\] put n=0,1,2,3,4,5 and get the solutions.

OpenStudy (jmartinez638):

That simultaneously makes more sense and is more complicated. Haha I like it

OpenStudy (anonymous):

correction it is \[z^6=64 e ^{\iota 2n \pi}\] then \[z=2e^\iota \frac{ 2n \pi }{ 3 }\]